OpenStudy (anonymous):

The sum of two numbers is 170 and their difference is 26. Find the two numbers. Write the solution as an ordered pair.

OpenStudy (michele_laino):

your problem can be modeled by the subsequent system: \[\Large \left\{ \begin{gathered} m + n = 170 \hfill \\ m - n = 26 \hfill \\ \end{gathered} \right.\] where m, and n are the requested numbers

OpenStudy (anonymous):

how do i put it in an order pair

OpenStudy (michele_laino):

you have to compute their values, first

OpenStudy (anonymous):

i don't know how to do that

OpenStudy (michele_laino):

for example, solve the second equation for m, what do you get?

OpenStudy (anonymous):

i don't know

OpenStudy (anonymous):

is it 72 and 98?

OpenStudy (michele_laino):

another way, is: if I sum side by side those 2 equations together, I get: \[\Large m + n + m - n = 170 + 26\] please simplify

OpenStudy (michele_laino):

add*

OpenStudy (mathmate):

Alternatively, in the \(special\) case when sum and difference are given, we can do the following: bigger number = (sum+difference)/2 smaller number = (sum-difference)/2 For example: Given sum = 20, difference = 8 bigger number = (20+8)/2 = 14 smaller number = (20-8)/2 = 6